# Fully Adaptive Zeroth-Order Method for Minimizing Functions with Compressible Gradients

### v2

@@ 5 Conclusion
In this paper, we presented a fully adaptive variant of ZORO, originally proposed in [6], for derivative-free minimization of a smooth functions with $n$ variables and $L$-Lipschitz continuous gradient. The new method, called ZORO-FA, does not require the knowledge of the Lispchitz constant $L$ or the effective sparsity level (when the gradients are compressible). At each iteration, ZORO-FA attempts to exploit a possible compressibility of gradient to compute a suitable gradient approximation using fewer than $\mathcal{O}\left(n\right)$ function evaluations. When the corresponding trial points fail to ensure a functional decrease of $\mathcal{O}\left(\epsilon^{2}\right)$, the method computes gradient approximations via forward finite-differences, which guarantee that the trial points will yield the desired functional decrease. With this safeguard procedure, we proved that, with probability one, ZORO-FA needs no more than $\mathcal{O}\left(n^{2}\epsilon^{-2}\right)$ function evaluations to find an $\epsilon$-approximate stationary point of the objective function. When the gradient vectors of the objective function are $p$-compressible for some $p\in(0,1)$, we proved that, with high probability, ZORO-FA has an improved worst-case complexity of $\mathcal{O}\left(s(\theta,p)\ln(n)\epsilon^{-2}\right)$ function evaluations, where $\theta\in(0,1/2)$ is a user-defined parameter that controls the relative error of the gradient approximations, and $s(\theta,p)$ is the effective sparsity level as defined in Corollary 2.6. Our preliminary numerical results indicate that ZORO-FA is able to exploit the presence of gradient compressibility, while being robust towards its absence.

### v3

@@ 5 Conclusion
In this paper, we presented a fully adaptive variant of ZORO, originally proposed in cai2022zeroth , for derivative-free minimization of a smooth functions with $n$ variables and $L$-Lipschitz continuous gradient. The new method, called ZORO-FA, does not require the knowledge of the Lipschitz constant $L$ or the effective sparsity level (when the gradients are compressible). At each iteration, ZORO-FA attempts to exploit a possible compressibility of gradient to compute a suitable gradient approximation using fewer than $\mathcal{O}\left(n\right)$ function evaluations. When the corresponding trial points fail to ensure a functional decrease of $\mathcal{O}\left(\epsilon^{2}\right)$, the method computes gradient approximations via forward finite-differences, which guarantee that the trial points will yield the desired functional decrease. With this safeguard procedure, we proved that, with probability one, ZORO-FA needs no more than $\mathcal{O}\left(n^{2}\epsilon^{-2}\right)$ function evaluations to find an $\epsilon$-approximate stationary point of the objective function. When the gradient vectors of the objective function are $p$-compressible for some $p\in(0,1)$, we proved that, with high probability, ZORO-FA has an improved worst-case complexity of $\mathcal{O}\left(s(\theta,p)\ln(n)\epsilon^{-2}\right)$ function evaluations, where $\theta\in(0,1/2)$ is a user-defined parameter that controls the relative error of the gradient approximations, and $s(\theta,p)$ is the effective sparsity level as defined in Corollary 1. Our preliminary numerical results indicate that ZORO-FA is able to exploit the presence of gradient compressibility, while being robust towards its absence.
Acknowledgements
We are very grateful to the associate editor and the two anonymous reviewers for their valuable comments and suggestions, which helped us improve this work.